CO/CO2+//Ni2+/Ni
Calculate the cell potential and energy
1
Expert's answer
2020-02-28T12:12:23-0500

"CO + H_2O \to CO_2 + 2H^+ + 2e^-" "E_o=0.53V"

"Ni^{2+} + 2e^- \to Ni" "E_o=-0.25V"

"\therefore E_{cell}=0.53-0.25=0.28V"


"\Delta G=-nEF=-2*0.08*96500=-15.44kJ" (Answer)